OpenStudy (anonymous):

Find all real solutions of the equation. (Enter your answers as a comma-separated list. If there is no real solution, enter NO REAL SOLUTION.) 2x + the square root of x + 13 = 10

OpenStudy (superdavesuper):

Start by moving the square root of x to one side of = n everything else to the other side...

OpenStudy (superdavesuper):

Okay I will help by moving...\[\sqrt{x} = 2x...\] then what?

OpenStudy (anonymous):

okay i got square root of x+13=10-2x

OpenStudy (superdavesuper):

Move the number to the left side plz? :)

OpenStudy (anonymous):

too which then I assume we must get ride of the square root so we square the entire thing giving us (square root of x=13)squared=(10-2x)squared or what is it that you are suggesting?

OpenStudy (superdavesuper):

:) u hardly need my help at all - yup dats da plan!

OpenStudy (superdavesuper):

oh wait wait u need to move 13 over the right first though

OpenStudy (anonymous):

(square root of x+13)squared my bad but then we get x+13=(10-2x)(10-2x) which then we foil the 10-2x side right?

OpenStudy (superdavesuper):

(square root of x +13) squared will give x + 2x13x square root of x so that is no good. U should move 13 over the right before squaring,

OpenStudy (anonymous):

so it's Square root of x=(10-2x-13)squared right?

OpenStudy (superdavesuper):

Simplify 10-2x-13 first...then square both side of the equation...

OpenStudy (anonymous):

\[(\sqrt{x+13})^{2}=(10-2x)^{2}\]

OpenStudy (superdavesuper):

Oh sorry I mis-read ur Q - didnt know the 13 is within the SQRT. So yes u are on the right track :)

OpenStudy (anonymous):

sorry first time using this place. So after all this we foil the (10-2x)(10-2x) we get \[x+13=100-20x-20x -4x^{2} \] which we combine the 20xs into 40x so we have once more \[x+13=\frac{ 100-40x+4x ^{2} }{ 4 }\]

OpenStudy (superdavesuper):

Now u need to move everything to one side n combine n simplify to a standard quadratic eqn: ax^2 + bx + C = 0 to solve.

OpenStudy (superdavesuper):

U are doing fine - sorry for my late replyl I was away...

OpenStudy (anonymous):

thanks man for you're help

OpenStudy (superdavesuper):

welcome u have helped urself the most :)